OpenStudy (rperry309):

PLEASE HELP Which graph represents the system 3x + 5y ≤ 9 and 2x + y ≥ 5?

OpenStudy (mathmale):

I'd suggest that YOU, r perry, actually graph these two inequalities: 3x + 5y ≤ 9 and 2x + y ≥ 5. Start by graphing 3x + 5y = 9 and 2x + y = 5, and then decide which side of each line should be shaded. How would you graph 3x + 5y = 9?

OpenStudy (mathmale):

Welcome to OpenStudy! My goal here is to guide you towards solving these problems yourself.